OpenStudy (anonymous):

The surface area of a cube is 600 cm2. What is the area of one face of this cube?

OpenStudy (anonymous):

S.A. of a cube = 6x^2 ; 6x^2 = 600 ; x=10 ; area of one face = 10 * 10 = 100

OpenStudy (anonymous):

i think it 100cm but want a second opinion

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

How many faces does a cube have? Think of a pair of dice. On one you have the numbers one to six. (That's a clue.) So, if 6 sides total 600 what does one side equal? Hint: Divide 600 by 6 to get your answer.